OpenStudy (anonymous):

) PLEASE HELP !!!!! You are asked to tutor a friend in algebra with adding and subtracting trinomials. Create an example of how to add two trinomials and how to subtract two trinomials. Come up with two trinomials and be sure to: a. Write out each step used to complete each addition and subtraction example. b. Show all the steps used to simplify the expressions. c. Provide the resulting degree and classification of each example. d. Explain why the polynomial is or is not closed over addition or subtraction based on the problem provided.

OpenStudy (anonymous):

do you still need help on this?

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

could you help me ?

OpenStudy (anonymous):

Addition: (3x^2+4x-2) + (-7x^2-10x+17) (add each part 3x^2 + -7x^2 = -4x^2 Then 4x + -10x = -6x Then -2+17= 15 So the answer to : (3x^2+4x-2) + (-7x^2-10x+17) is... -4x^2 - 6x + 15 (this is a second degree trinomial)

OpenStudy (anonymous):

Subtraction: (3x^2+4x-2) - (-7x^2-10x+17) (subtract each part) - like terms 3x^2 - -7x^2= 10x^2 4x - -10x= 14x -2-17= -19 So the answer to (3x^2+4x-2) - (-7x^2-10x+17) is.... 10x^2 + 14x - 19x (second degree trinomial)

OpenStudy (anonymous):

But im not really sure what d is asking...?

OpenStudy (anonymous):

ok, thank you so much... & I'm not sure what its asking either.

OpenStudy (broskishelleh):

@ImaGuest Do you need this?

OpenStudy (anonymous):

@Broskishelleh No, I don't

OpenStudy (anonymous):

Sorry, I accidentaly opened this page looking for something else

OpenStudy (anonymous):

Didn't realize I still had it open

OpenStudy (broskishelleh):

Oh, ok do you want to go back to your other question

OpenStudy (anonymous):

yes